• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name dalton

    Log entry time 00:26:31 on December 14, 2009

    Entry number 306748

    This entry is a followup to: 306732

    keyword=Compton problems continue

    We have not made progress in making the laser lock. We tried a power cycle of the locking elctronics, a couple of epics reboots and a compton computer reboot but to no avail.

    The symptoms are kind of bizarre. The system appears to be flashing but the camera on the table seems to have stopped working so we can't tell what mode it is. The reflection dip is deep (~ 50% somtimes) so it's probably the 00 mode.

    When the servo is on and the Tracking Level is set at 0.65 (where we have had it throughout HAPPEx and before) then the temperature ramp continues as though it were unlocked while the slow control jumps around to crazy values. The slow control should only change when the system is locking. So there some internal confusion as to whether it is trying to lock or to scan. This behaviour occurs for Tracking Levels between 0.61 and 0.66. The slow control steers the laser leading to a saw tooth pattern in the photodiode positions. The positions themselvs are pretty much what they were when the system last worked so alignment is not suspected.

    We're at a loss as to what to try now. Any ideas would be appreciated.

    A copy of this log entry has been emailed to: nanda@jlab.org, rahim@jlab.org, camsonne@jlab.org, mfriend@jlab.org,dalton@jlab.org, paschke@jlab.org,xiaochao@jlab.org,rom@jlab.org,reimer@jlab.org,dwang@jlab.org



    Figure 1



    Figure 2